A transformer has 1000 turns on its primary side which uses a 120 V AC source. If the transformer is used to produce 10 W of power at 12.0 V on the secondary side, and losses are negligible, determine:

a) the number of turns on the secondary side. b) the current on the secondary side.


1
Expert's answer
2021-06-22T13:33:30-0400

(a) We can find the number of turns on the secondary side as follows:


"\dfrac{V_2}{V_1}=\dfrac{N_2}{N_1},""N_2=\dfrac{V_2}{V_1}N_1=\dfrac{12\ V}{120\ V}\cdot1000\ turns=100\ turns."

(b) We can find the current on the secondary side from the definition of the power:


"P_2=I_2V_2,""I_2=\dfrac{P_2}{V_2}=\dfrac{10\ W}{12\ V}=0.83\ A."